SIPB Services

The most essential inessential computing at MIT

SIPB services are free and available to all members of the MIT community. Below are some of our most popular services.

• scripts.mit.edu scripts.mit.edu is a service that allows you to host web-accessible CGI scripts out of your Athena locker (there is no official IS&T service that lets students host dynamic content). Perl, PHP, Python, and many other languages are supported, and we offer free MySQL databases as well. Easy auto-installers also let you set up a blog, wiki, or other software for yourself or a student group with a single command on Athena. scripts.mit.edu also offers a mail script service and a cron script service. See http://scripts.mit.edu for more information.

• linerva.mit.edu Linerva is SIPB’s Linux dialup service. The official Athena dialup service, athena.dialup.mit.edu, runs Solaris (and thus lacks a lot of useful Linux-only software), is slow, and logs users out after a few hours. Linerva is our answer to these problems, with a number of other advantages. See http://linerva.mit.edu for more information.

• Debathena Debathena brings the full functionality of an Athena workstation to your Debian or Ubuntu system through a set of Debian packages. You can easily install, and cleanly uninstall, whatever packages you want—from a few of the client utilities, to a full-fledged Kerberos login system supporting AFS home directories and locker software. Debathena is also the base for Linerva, the SIPB public Linux dialup. See http://debathena.mit.edu for more information.

• Library Access to Music Project LAMP, the Library Access to Music Project, is an electronic music library containing 1840 CDs, where you can request songs at http://lamp.mit.edu and they will be played on MIT cable TV.

• The SIPB AFS Cell and Locker Software SIPB maintains hundreds of useful pieces of software in AFS lockers that are not installed locally on Athena. While many popular programs (like the mutt email client, various compilers, and the imapback email backup script) are available in the sipb and outland lockers, many other programs are located in their own lockers in SIPB’s AFS cell. SIPB maintains the whichlocker program in the outland locker, which makes it easy to find locker software.

• IAP Classes and Cluedumps SIPB organizes dozens of intensive classes on a variety of computing topics during IAP. Additionally, SIPB offers a series of Cluedump talks during the Fall term, with a different topic every week. Add yourself to cluedump-announce@mit.edu if you’re interested.

• Your Project Here SIPB can offer resources and help for making your project to improve MIT computing happen!